# Connecting the Dots: UV-Bright Companions of Little Red Dots as Lyman-Werner Sources Enabling Direct Collapse Black Hole Formation

### v1

@@ Abstract
We compile a sample of 83 Little Red Dots (LRDs) with JWST imaging and find that a substantial fraction ($\sim$43%, rising to $\gtrsim$85% for the most luminous LRDs) host one or more spatially offset, UV-bright companions at projected separations of $0.5$ kpc $\lesssim d\lesssim 5$ kpc, with median $\langle d\rangle=1.0$ kpc. This fraction is even higher when smaller spatial scales are probed at high S/N ratio: we show that the two most strongly lensed LRDs known to date, A383-LRD and the newly discovered A68-LRD, both have UV-bright companions at separations of only $d\sim 0.3$ kpc, below the resolution limit of most unlensed JWST samples. We explore whether these ubiquitous red / blue configurations may be physically linked to the formation of LRDs, in analogy with the “synchronized pair” scenario originally proposed for direct-collapse black hole formation. In this picture, ultraviolet radiation from the companions, which typically have modest stellar masses ($M_{*}\sim 10^{8-9}{\rm M_{\odot}}$), suppresses molecular hydrogen cooling in nearby gas, allowing nearly isothermal collapse and the formation of extremely compact objects, such as massive black holes or quasi-stars. Using component-resolved photometry and SED modeling, we infer Lyman–Werner radiation fields of $J_{21,\rm LW}\sim 10^{2.5}$$10^{5}$ at the locations of the red components, comparable to those required in direct-collapse models, suggesting that the necessary photodissociation conditions are realized in many LRD systems. This framework provides a simple and self-consistent explanation for the extreme compactness and distinctive spectral properties of LRDs, and links long-standing theoretical models for early compact object formation directly to a population now observed with JWST in the early universe.
@@ I Introduction
Interestingly, theoretical models have long predicted that the formation of extremely compact massive objects in the early Universe requires precisely such environments. In nearly metal-free gas, molecular hydrogen provides the dominant cooling channel, but far-ultraviolet photons in the Lyman-Werner (LW) band can dissociate H2 via the Solomon process (Draine and Bertoldi, 1996). If the local LW radiation field is sufficiently intense, molecular hydrogen is dissociated, and cooling is suppressed, allowing gas clouds to bypass normal star formation and collapse rapidly at the atomic-cooling threshold (Omukai, 2001), avoiding fragmentation (Oh and Haiman, 2002) and ultimately leading to a massive black hole (Bromm and Loeb, 2003; Volonteri and Rees, 2005). This has been envisioned to occur through the intermediate formation of a supermassive star (Agarwal et al., 2012; Latif et al., 2013; Ferrara et al., 2014; Sugimura et al., 2014; Hosokawa et al., 2016; Chon et al., 2016; Hirano et al., 2017; Haemmerlé et al., 2018), a short-lived quasi-star phase (e.g. Begelman et al., 2006; Volonteri and Begelman, 2010), or an extremely dense star cluster (e.g. Omukai et al., 2008; Devecchi and Volonteri, 2009; Boekholt et al., 2018; Natarajan, 2011) These pathways can all ultimately seed a massive ’heavy seed’ black hole (see Volonteri, 2010; Haiman, 2013; Inayoshi et al., 2020, for comprehensive reviews), but they rely on the suppression of efficient $\rm H_{2}$} cooling and fragmentation of a pristine cloud.
@@ VI Conclusions
We thank Earl Bellinger, Fabio Pacucci, Andrea Ferrara and Dale Kocevski for useful discussions. This work used the following python packages: Matplotlib (Hunter, 2007), SciPy (Virtanen et al., 2020), NumPy (van der Walt et al., 2011), AstroPy (Astropy Collaboration et al., 2022) and colossus (Diemer, 2018) and photutils (Bradley et al., 2025).
@@ Appendix A A strongly lensed Little Red Dot behind Abell 68
Figure 6: RGB composite of the Abell 68 cluster field, constructed from F [^4] W+F356W (red), F277W+F200W (green), and F150W+F115W+F090W (blue), showing the three lensed images of A68-LRD. The imaging is publicly available from the VENUS collaboration (Program ID: GO 6882; PI: S. Fujimoto).

### v2

@@ Abstract
We compile a sample of 83 Little Red Dots (LRDs) with JWST imaging and find that a substantial fraction ($\sim$43%, rising to $\gtrsim$85% for the most luminous LRDs) host one or more spatially offset, UV-bright companions at projected separations of $0.5$ kpc $\lesssim d\lesssim 5$ kpc, with median $\langle d\rangle=1.0$ kpc. This fraction is even higher when smaller spatial scales are probed at high S/N ratio: the two most strongly lensed LRDs known to date, A383-LRD and newly discovered A68-LRD, both have UV-bright companions at separations of only $d\sim 0.3$ kpc, below the resolution limit of most unlensed JWST samples. We explore whether these ubiquitous red / blue configurations are physically linked to the formation of LRDs, in analogy with the “synchronized pair” scenario originally proposed for direct-collapse black hole formation. In this picture, ultraviolet radiation from the companions, with typically modest stellar masses ($M_{*}\sim 10^{8-9}{\rm M_{\odot}}$), suppresses molecular hydrogen cooling in nearby gas, allowing nearly isothermal collapse and the formation of extremely compact objects, such as massive black holes or quasi-stars. Using component-resolved photometry and SED modeling, we infer Lyman–Werner radiation fields of $J_{21,\rm LW}\sim 10^{2.5}$$10^{5}$ at the locations of the red components, comparable to those required in direct-collapse models, suggesting the necessary photodissociation conditions are realized in many LRD systems. This framework provides a simple and self-consistent explanation for the extreme compactness and distinctive spectral properties of LRDs, and links long-standing theoretical models for early compact object formation directly to a population now observed with JWST in the early universe.
@@ I Introduction
Interestingly, theoretical models have long predicted that the formation of extremely compact massive objects in the early Universe requires precisely such environments. In nearly metal-free gas, molecular hydrogen provides the dominant cooling channel, but far-ultraviolet photons in the Lyman-Werner (LW) band can dissociate H2 via the Solomon process (Draine and Bertoldi, 1996). If the local LW radiation field is sufficiently intense, molecular hydrogen is dissociated, and cooling is suppressed, allowing gas clouds to bypass normal star formation and collapse rapidly at the atomic-cooling threshold (Omukai, 2001), avoiding fragmentation (Oh and Haiman, 2002) and ultimately leading to a massive black hole (Loeb and Rasio, 1994; Bromm and Loeb, 2003; Volonteri and Rees, 2005; Lodato and Natarajan, 2006, 2007). This has been envisioned to occur through the intermediate formation of a supermassive star (Agarwal et al., 2012; Latif et al., 2013; Ferrara et al., 2014; Sugimura et al., 2014; Hosokawa et al., 2016; Chon et al., 2016; Hirano et al., 2017; Haemmerlé et al., 2018), a short-lived quasi-star phase (e.g. Begelman et al., 2006; Volonteri and Begelman, 2010), or an extremely dense star cluster (e.g. Omukai et al., 2008; Devecchi and Volonteri, 2009; Boekholt et al., 2018; Natarajan, 2011) These pathways can all ultimately seed a massive ’heavy seed’ black hole (see Volonteri, 2010; Haiman, 2013; Inayoshi et al., 2020, for comprehensive reviews), but they rely on the suppression of efficient $\rm H_{2}$} cooling and fragmentation of a pristine cloud.
@@ VI Conclusions
We thank Earl Bellinger, Fabio Pacucci, Andrea Ferrara and Dale Kocevski for useful discussions. This work is based on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with programs 1345, 1180, 1181, 1286, 1243, 6882, 2561, 1324, 4111, 1895. The Cosmic Dawn Center (DAWN) is funded by the Danish National Research Foundation under grant DNRF140. This work used the following python packages: Matplotlib (Hunter, 2007), SciPy (Virtanen et al., 2020), NumPy (van der Walt et al., 2011), AstroPy (Astropy Collaboration et al., 2022) and colossus (Diemer, 2018) and photutils (Bradley et al., 2025).
@@ Appendix A A strongly lensed Little Red Dot behind Abell 68
Figure 6: RGB composite of the Abell 68 cluster field, constructed from F [^4] W (red), F277W (green), and F150W (blue), showing the three lensed images of A68-LRD. The imaging is publicly available from the VENUS collaboration (Program ID: GO 6882; PI: S. Fujimoto).